Crew 218 Journalist Report 29Dec2019

[title Journalist Report – December 29th]

[category journalist-report]

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 07

"Hydrate or die-drate!"

This was our mantra of the day as we prepared for our first EVA in a week. There’s a knife’s edge to tiptoe between under-insulating and over-bundling before heading out onto the snowy Martian terrain. The consequences of playing your cards wrong are either suffering icy numbness or heat exhaustion. Regardless of your position on the thermal spectrum, lugging around the EMU (Extravehicular Mobility Unit [my favorite acronym]) is some serious cardio. The only thing you’ll be burning more than calories is precious H2O, and you can’t drink while in EMU mode, so drink up beforehand and maybe store some in your cheeks too. But don’t forget that our EMUs don’t have lavatory functionality yet, so don’t drink too much. Basically, if your EVA is longer than two hours, godspeed.

Knowing that we were rusty with EVA procedures, we began preparing for our expedition over an hour prior. I reassembled the antenna I built for my research project and gave it one last test before its maiden voyage. Antenna is a strong word – it’s more like a haphazard amalgamation of PVC, copper wire, and coax cable. In the hopes that the electrical tape would hold together, we began suiting up. Pat, Jonathan, and Shefali helped LuzMa, Cesare, and I wriggle into our suits. We meticulously checked and double-checked every aspect of our expeditionary ensembles. Hats, gloves, bandannas, radios, microphones, EMUs on, straps tightened, helmets secured – ready to rock and roll.

Cue the fog machine and dramatic music; we stepped into the airlock. Pause the music, please allow five minutes for full depressurization. The time drags on as the 3 of us are sardine’d in the airlock cylinder with my PVC Frankenstein’s monster. Four minutes remaining. Three. As the air gets thinner our suits get tighter, something seems amiss. Why do my feet feel so… comfortable? The tape securing my microphone to my face contorts as my smile vanishes. I pretzel my spine to get downward view with the helmet on and sure enough: I’m still wearing my slippers. All of that time preparing for our first EVA in a while, and I overlooked my footwear of all things. Here we are about to embark on a monumental exploratory enterprise and I’m equipped like a middle-aged man stepping out to fetch this morning’s copy of The Martian Times.

I immediately alerted the crew and we began repressurizing the chamber. The next few minutes were slated to be agony as I had to stand there and receive infinite (deserved) ridicule over the radio. However, as if it were destiny, LuzMa’s hat jumped off of her head within the EMU helmet. She squirmed her head around the expensive fishbowl trying to reclaim her beanie, but to no avail. I’m glad the radio microphones are muted by default, because I was cackling uncontrollably in my helmet. I now stand by the claim that my shoe mishap was a noble sacrifice of dignity to rescue LuzMa from her impending headwear disaster. I’m never gonna live that down.

We re-assimilated with the Habitat air and fixed all of our wardrobe malfunctions. That was the practice round; now for the real deal. We must have shaken off all the rust on the first attempt, because our second attempt went flawlessly! We ventured out into the frozen wastes and returned to Marble Ritual, the same destination as our training EVA so many moons ago. But this time we were heading there to gather data for the first time. There was something sentimental about returning to that site one week later as full-fledged Martian survivors and scientists. We gathered geological, meteorological, and radio frequency data for our respective personal projects and returned to the Hab safe and sound (and sweaty).

We slithered out of our EMU suits accomplished and exasperated, wrapping up our EVA. With that excitement at its end, we immersed ourselves in data parsing and the energy of the day quickly waned. Before we knew it, the sun and moon had traded places and it was time for reports and dinner. I write this now through watery eyes from accidentally inhaling cayenne pepper powder. I think it’s time for me to concede that today is not my day and retire for the night. It’s bright and early tomorrow for some more potential EVAs. We’ve got a week left to make up for lost science time, let’s make the best of it!

Commander Report – December 29th

Crew 218 Commander Report 29-12-2019

Sol 7 – Half-way

We are in sol 7 of 13, reaching half-way through our mission. The day was extremely cold, all the pipes froze, and we have not been able to pump water up in the habitat, so we are working (really well) with rationed water. I began recording my biometrics for one of the research projects. Finally, three of us managed to get outside in a short EVA to continue two research projects. The landscape is as amazing as ever, and white is giving way to a great Martian red. The afternoon was spent between research and preparation of samosa, and the crew keeps getting the best of all the situations. Mood is great and we are ready to spend some more time outside.

Crew 218 Sol Summary 29Dec2019

[title Sol Summary – December 29th]

[category sol-summary]

Sol: 07

Summary Title: Breath of Fresh Martian Air

Author’s name: Pat Pesa

Mission Status: Crew is excited about Today’s progress and the clear weather ahead.

Sol Activity Summary: Crew 218 Finally got out of the Habitat for an EVA this afternoon! Along with continued work on our Indoor/Outdoor Climate monitoring, we started our Journalist’s project in Detecting Radio Signal Strength in the local area.

Look ahead plan: Hoping for more EVAs for tomorrow after more melting occurs. Tomorrow we will also try to finish crew interviews and training videos.

Anomalies in work: None

Weather: Partly Cloudy, around 25 degrees

Crew physical status: Healthy

EVA: none

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Crew 218 Science Report 29Dec2019

[title Science Report – December 29th]

[category science-report]

Crew 218 Science Report 29-DEC-19
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

No EVA: nothing to report.

3. Subsurface structure on Mars:

No EVA: nothing to report.

4. Detecting radio signal strength:

No EVA: nothing to report.

5. EVA workload analysis:

No EVA: nothing to report.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

No EVA: nothing to report.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Nothing to report.

8. Messier and other space objects for outreach:

Cloudy Weather: nothing to report.

9. Reliability and maintenance:

No EVA: nothing to report.

10. Medical readings in preparation for future crew-wide project:

The Commander, Dr. Cesare Guariniello, has donned the Zephyr bio-medical sensor at 11:00AM, and is testing real-time measurement infrastructure for EVA and habitat medical monitoring. He will wear the instrument continuously for 24-48 hours.

11. Collection of clay, shale, and hematite samples:

No EVA: nothing to report.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

Crew 218 GreenHab Report 29Dec2019

[title GreenHab Report – December 29th]

[category greenhab-report]

Crew 218 GreenHab Report 29-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23°C; 18%

11:00AM

Floor Unit: 18°C

Electronic: 24°C

humidity 18%

12:35PM

Floor Unit: 18°C

Electronic: 23.7°C

humidity 18%

3:15PM

Floor Unit: 17°C

Electronic: 19.8°C

humidity 20%

5:45PM

Floor Unit: 14°C

Electronic: 19.2°C

humidity 20%

Max: 28.8; 21%

Min: 17.2°C; 16%

Hours of supplemental light: Light system 6:15PM-11:15AM.

Daily water usage for crops: 5.4Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~100 Gal.

Time(s) of watering for crops:

6:00PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: Zucchini, Cherry Tomato, and Pepper had first sprouts.

—Door was opened while in GreenHab, but closed upon leaving.

Harvest:

Aloe Vera 9g

Crew 218 EVA Report 29Dec2019

[title EVA Report – December 29th]

[category eva-report]

EVA #: 3

Author: Cesare Guariniello

Purpose of EVA: weather observation project; radio mapping project; collection of geological samples

Start time: 12:00

End time: 13:18 (including engineering EVA)

Narrative: one walking EVA has been approved for today. Due to the presence of snow, the commander joined the EVA for better knowledge of the terrain. Working on snow was easy and safe, and a couple of narrow streams formed by the melted water were less than one step across. At the end of the simulated depressurization cycle, the crew noticed that the front door was still blocked by some snow and Pat completed the shoveling process. At the same time, both Ben and Luz Ma experienced problems with their equipment, therefore a cycle of re-pressurization was initiated and the crew re-entered the habitat to fix a hat and a pair of shoes. After the second depressurization, Cesare, Luz Ma and Ben egressed and Luz Ma and Ben became their measurements with the help of Cesare. Then, the crew grabbed their bag with tools, EVA first aid kit, map, and rock hammer and walked towards the MDRS sign and out along Cow Dung Road to Marble Ritual. Radio measurements were collected along the way. Once at destination, both weather data and radio measurements were collected before the crew headed back to the habitat. An engineering EVA was performed before ingress.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): E518700, N4250800

EVA Participants: LuzMa Agudelo, ENG; Ben Durkee, JOU; Cesare Guariniello, CMD.

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Walking

Crew 218 Operations Report 29Dec2019

[title Operations Report – December 29th]

[category operations-report]

Crew 218 Operations Report 29-Dec-19

SOL: 07
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 11
From what time last night: 2000
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night: 78%
Diesel Reading – 77%
Station Propane Reading – 71%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 21
Water Meter (units): 0147111.1
Water (static tank) (gal): 395
Static to Loft Pump used – No
Water in Green Hab (gal): 100
Water in Science Dome (gal): 0
Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes
Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? No
CrewCar used and why, where? No
General notes and comments: N/A
Summary of the internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: The pipe from the static tank through the loft tank froze before we did the regular daily filling yesterday (12/28/2019). We tried to fill the tank several times during the day today without luck. The pipes are still frozen, we are rationing the water, and we will keep checking on the status of the pipe.

Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns, and requests to Mission Support: NA